# Understanding the differences in Foundation Models: Attention, State Space Models, and Recurrent Neural Networks

### v2

@@ I.2 WikiText-103 experiments
  • Optimizer and schedule: Weight decay of 0.1, linear warmup with duration of 10%, AdamW optimizer [Loshchilov and Hutter, 2019] with $\beta=(0.9,0.95)$, and gradient clipping $=1$. For each run, we sweep the learning rates in $[0.0003,0.001,0.003,0.01]$ and train for 50 epochs.
@@
  • Position information: Positional embeddings [Brown et al., 2020] are used as in standard Pythia.
@@
Appendix J Computational Resources
All experiments (MQAR and WikiText-103) were run on a cluster with $11$ nodes with the following GPU and CPU specifications:
@@
The MQAR training and test runs were parallelized and assigned to the best available GPU node, while the parallelized training on WikiText-103 was exclusively run on the NVIDIA A100 (80GB) node.
For each learning rate sweep of the MQAR runs described in Appendix I we estimate the average runtime to be 1h, [^10] Obviously, larger model sizes and MQAR tasks with larger sequence length took longer than smaller models and tasks with shorter sequence length. However, a more accurate time estimate is hard to obtain due to the cluster setup with multiple different GPU models and the fact that we terminated tasks early if the $99\%$ accuracy threshold was achieved. leading to a total unparallelized runtime of 54 days for all MQAR tasks. There where approximately $20$ runs for debugging and training purposes, which were terminated after a few minutes, thus we did not include them in the time estimate.
@@
Appendix K Extended Results on MQAR
In Figure 5 we report the complete results of all MQAR experiments detailed in Appendix I. A selected subset of these are already presented in Figure 1 and Figure 2 in the main text.
@@
While Figure 2 shows that normalized attention outperforms standard linear attention [Katharopoulos et al., 2020], Figure 5 shows an even more significant performance gap. Additionally, we note that SSD slightly outperforms S6, which was already hinted at in [Dao and Gu, 2024], and that normalized attention performs on par with S6. Together these results hint at the importance of normalization both for attention and SSMs. The comparison of S6 and SSD shows that reducing the number of parameters in the state transition $\Lambda_{i}$ from $d$ to a scalar does not hurt performance, which is further supported by the findings in [Dao and Gu, 2024]. These experiments also suggest that the recursive structure in $\Lambda_{i}$ and $B_{i}$ (present in S6 and SSD but not in normalized attention or linear attention) is less important than proper normalization of the attention scores. Additionally, the results on WikiText-103 (Table 1) show that better normalization can close 25% of the perplexity gap between linear attention and softmax attention. Together, these results warrant more investigations into new and better normalization techniques for attention-based models.

### v3

@@ Appendix I S6 uses reversed sigmoid in state transition matrix
  • Optimizer and schedule: Weight decay of 0.1, linear warmup with duration of 10%, AdamW optimizer [59] with $\beta=(0.9,0.95)$, and gradient clipping $=1$. For each run, we sweep the learning rates in $[0.0003,0.001,0.003,0.01]$ and train for 50 epochs.
@@
  • Position information: Positional embeddings [60] are used as in standard Pythia.
@@
Appendix K Computational Resources
All experiments (MQAR, LRA, and WikiText-103) were run on a cluster with $11$ nodes with the following GPU and CPU specifications:
@@
The MQAR and LRA training and test runs were parallelized and assigned to the best available GPU node, while the parallelized training on WikiText-103 was exclusively run on the NVIDIA A100 (80GB) node.
For each learning rate sweep of the MQAR runs described in Appendix J we estimate the average runtime to be 1h, [^15] Obviously, larger model sizes and MQAR tasks with larger sequence length took longer than smaller models and tasks with shorter sequence length. However, a more accurate time estimate is hard to obtain due to the cluster setup with multiple different GPU models and the fact that we terminated tasks early if the $99\%$ accuracy threshold was achieved. leading to a total unparallelized runtime of 54 days for all MQAR tasks. There where approximately $20$ runs for debugging and training purposes, which were terminated after a few minutes, thus we did not include them in the time estimate.
For each task of the LRA benchmark, we estimate the average runtime to be 2h for Image, 5h for Text, 6h for ListOps, 30h for Retrieval, and 45h for Pathfinder, leading to a total unparallelized runtime of 31 days for all LRA tasks. Note that to obtain better hyperparameters, they would need to be tuned for each task, which would significantly increase the total runtime. There where approximately $30$ runs for debugging and training purposes, which were terminated after a few minutes, thus we did not include them in the time estimate.
@@
Appendix L Extended Results on MQAR
Figures 5 and 6 show the MQAR results of Figures 2 and 3 but for multiple runs over 10 different seeds.
[…]
Figure 6: Comparison of qLSTM (8) and a qLSTM variant where the original state transition $\Lambda_{i}$ is replace by (23). Solid lines are the average accuracy over 10 different seeds, while the shaded area show the standard deviation.
In Figure 7 we report the complete results of all MQAR experiments detailed in Appendix J. A selected subset of these are already presented in Figure 1 and Figure 2 in the main text.
@@
While Figure 2 shows that normalized attention outperforms standard linear attention [14], Figure 7 shows an even more significant performance gap. Additionally, we note that SSD outperforms S6, which was already hinted at in [8], and that normalized attention performs on par with S6. Together these results hint at the importance of normalization both for attention and SSMs. The comparison of S6 and SSD shows that reducing the number of parameters in the state transition $\Lambda_{i}$ from $d$ to a scalar does not hurt performance, which is further supported by the findings in [8]. These experiments also suggest that the recursive structure in $\Lambda_{i}$ and $B_{i}$ (present in S6 and SSD but not in normalized attention or linear attention) is less important than proper normalization of the attention scores. Additionally, the results on WikiText-103 (Table 1) show that better normalization can close 25% of the perplexity gap between linear attention and softmax attention. Together, these results warrant more investigations into new and better normalization techniques for attention-based models.
@@
Appendix M Extended Results on LRA
In Table 3 we report the complete results of all LRA experiments detailed in Appendix J. The average performance over all tasks is reported in Table 1 in the main text.
While normalized attention (22) performs slightly better on LRA than the other two attention-based models, there is a significant performance gap to the S6 model (a selective SSM variant). The reason for this gap potentially lies in the recurrent normalization employed in selective SSMs (see Section 4.2). Interestingly, S6 only achieves significantly higher accuracy on the tasks Text and Image, showing that selective SSM models are particularly suited for long-range classification of text and image modalities.
Table 3: Model performance in terms of test accuracy on the LRA benchmark. The first entry (Random) represents the performance of random guessing on the task, i.e., indicating the baseline above which a model is considered to have learned a meaningful representation.
Model LRA Task [%]
ListOps Text Retrieval Image Pathfinder avg
Random 10.00 50.00 50.00 10.00 50.00 34.00
Softmax Attention [3] 35.72 63.10 77.46 34.22 69.32 55.96
Linear Attention [14] 16.12 64.41 76.44 37.97 72.64 53.52
Normalized Attention (22) 38.24 64.96 79.68 35.96 71.56 58.08
S6 [7] 38.02 81.34 80.50 65.08 69.26 66.84